OpenStudy (mackenzie2013):

Find the surface area of a triangular prism that has the following dimensions. Use the formula SA = 2B + Ph. B = 84 mm2 P = 56 mm h = 5 mm

OpenStudy (mackenzie2013):

@ganeshie8 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Plug in the values to get... SA = 2B + Ph SA = 2*84+56*5 SA = 448 The surface area is 448 square mm. Hopefully this is making sense.